The calorie - nutrient information for an apple and an avocado is given in the table. How many of each should be eaten to get exactly 1000 calories and 100 grams of carbohydrates?

One Apple One Avocado
Calories 100 350
Carbohydrates (grams) 24 14
________________________________________

Answer by mananth(16946)


let apples be x
avocado be y
100x +350y=1000
divide by 50
2x+7y=20
multiply equation by -2
-4x-14y=-40
24x+14y=200
Add the equations
20x=160
x= 8 apples
plug value of x in first equation
16+7y=20
7y=4
y=4/7 avocado